OpenStudy (anonymous):

compute the linear equation (-22.2 , 19.9)+3(-0.8 , 6.3)

OpenStudy (dan815):

?

OpenStudy (dan815):

Could you clarify a little more? Do they want us to write a linear equation that corresponds to this vector?

OpenStudy (anonymous):

It says compute the following linear combinations, I think it means to solve it. The first one was 4(1,7)-2(5,-3) and i got 64

OpenStudy (dan815):

Okay hold on I know someone who would know, @ikram002p

OpenStudy (ikram002p):

@brookdrew how did u get 64 for second eq ?

OpenStudy (anonymous):

I did 4(1,7)-2(2) then 4*(1,7)-2(2) then 68-2(2) then 68-4

OpenStudy (ikram002p):

well no its not like that

OpenStudy (ikram002p):

u must get a point as an answer

OpenStudy (anonymous):

oh okay, how is it then ?

OpenStudy (ikram002p):

u have :- 4(1,7)-2(5,-3) = (1*4, 7*4)-(2*5, 2*-3)=(4,28)-(10,-6)=(4-10 , 28+6)=(-6,34)

OpenStudy (ikram002p):

so can u solve this the same way ? (-22.2 , 19.9)+3(-0.8 , 6.3)

OpenStudy (anonymous):

so (-22.2, 19.9) (3*-0.8. 3*6.3) ?

OpenStudy (anonymous):

if its right il go to the next step ?

OpenStudy (ikram002p):

dnt forget there is + sighn btw

OpenStudy (anonymous):

oh okay, so is the rest right ?

OpenStudy (ikram002p):

(-22.2, 19.9)+ (3*-0.8. 3*6.3) so compute 3*-0.8 =? and 3*6.3 = ?

OpenStudy (anonymous):

(2.4,18.9)

OpenStudy (anonymous):

Is the final answer (-19.8,38.8)

OpenStudy (ikram002p):

(-2.4, 18.9) thats correct nw compute (-22.2, 19.9)+ (-2.4, 18.9) =?

OpenStudy (anonymous):

(-19.8,38.8)

OpenStudy (ikram002p):

-22.2+ - 2.4 =?? its not -19.8

OpenStudy (anonymous):

so (-24.6,38.8)

OpenStudy (anonymous):

-22.2+-2.4=-24.6

OpenStudy (ikram002p):

yep , correct ^_^

OpenStudy (anonymous):

Thank youu!!

OpenStudy (ikram002p):

np :)

OpenStudy (anonymous):

:)